# Minimum Weighted Feedback Arc Sets for Ranking from Pairwise Comparisons

### v1

@@ Abstract
The Minimum Weighted Feedback Arc Set (MWFAS) problem is fundamentally connected to the Ranking Problem - the task of deriving global rankings from pairwise comparisons. Recent work [He et al. ICML2022] has advanced the state-of-the-art for the Ranking Problem using learning-based methods, improving upon multiple previous approaches. However, the connection to MWFAS remains underexplored. This paper investigates this relationship and presents efficient combinatorial algorithms for solving MWFAS, thus addressing the Ranking Problem. Our experimental results demonstrate that these simple, learning-free algorithms not only significantly outperform learning-based methods in terms of speed but also generally achieve superior ranking accuracy. [^1] This paper is a preliminary report.
@@ 1.2.1 Previous Work About Minimum Weighted Feedback Arc Set
  • The Minimum Weighted Feedback Arc Set (MWFAS) problem has received significant attention, particularly regarding approximation algorithms. Early work by [10] provided an approximation algorithm for the MWFAS problem, followed by improvements by [9], that introduced a divide-and-conquer approach using spreading metrics. Although effective, these algorithms used the ellipsoid method for solving linear programming relaxations, making them computationally slow in practice.
@@
    [19] proposed an $O(VE^{4})$ algorithm for MWFAS. However, as the number of edges in the graph increases, the algorithm’s performance becomes slower. Our work implemented some of these algorithms, and the results, available in our GitHub repository, confirm the practical limitations regarding their speed.
    [26] addresses the parameterized complexity of the feedback vertex and arc set problems in tournaments, focusing on both weighted and unweighted versions. It presents fixed parameter tractable (FPT) algorithms for these problems, including an $O((2.4143)^{k}n^{\omega})$ algorithm for the weighted feedback vertex set problem and an $O((2.415)^{k}n^{\omega})$ algorithm for the weighted feedback arc set problem where $\omega$ is exponent of the running time of the best matrix multiplication algorithm, and $k$ is the size of the feedback set. The paper also introduces algorithms for the parametric duals of these problems and explores their complexity in oriented directed graphs and general directed graphs.
    [17] [28] explores the use of satisfiability-modulo theories (SMT) solvers to solve the Minimum Feedback Arc Set (MFAS) problem. It presents an SMT encoding for MFAS and highlights the use of incremental solving features, such as unsatisfiable cores. The results show that while the SMT-based approach has potential, it remains computationally infeasible for larger instances, with solving times that scale poorly. The authors also provide complexity bounds, demonstrating that the MFAS problem remains challenging even with advanced SMT techniques.
@@ 1.2.2 Previous Work About Ranking Problem
The ranking problem has also been studied extensively in the literature. [7] introduced Sync-Rank, a method that employs eigenvector and semidefinite programming synchronization to solve ranking, constrained ranking, and rank aggregation problems. The approach leverages spectral or SDP relaxations, followed by rounding procedures to recover rankings. Sync-Rank outperformed other algorithms in both synthetic and real-world datasets while supporting semi-supervised ranking and identifying consistent partial rankings. [6] addressed the problem of ranking players using incomplete pairwise comparison data within the Bradley-Terry-Luce model framework. It established the minimax rate for this ranking task with respect to Kendall’s tau distance, identifying a unique transition between exponential and polynomial rates based on the signal-to-noise ratio. The paper proposed a divide-and-conquer algorithm that achieved the minimax rate, demonstrating its theoretical optimality and computational efficiency.
@@
[15] introduces a novel pipeline called Graph Neural Re-Ranking (GNRR), which integrates Graph Neural Networks (GNNs) into the re-ranking phase of information retrieval systems. By constructing corpus graphs to model semantic and lexical document relationships, the GNRR approach captures cross-document interactions, improving ranking performance on popular metrics compared to traditional uni-variate baselines. Experimental results show consistent performance gains, especially in challenging datasets, highlighting the potential of GNNs to refine document rankings effectively. The most similar work to ours is [5], which focuses on vertex ordering in directed graph streams. They address the Feedback Arc Set problem and propose removing feedback arcs for conducting topological sorts. However, unlike our work, their focus is on unweighted tournament graphs without experimental validation, whereas we address weighted general graphs with extensive experiments.
[4] presents exact algorithms for solving several vertex ordering problems on graphs. The proposed algorithms operate in either $O^{*}(2^{n})$ time and space or $O^{*}(4^{n})$ time and polynomial space, leveraging dynamic programming and recursive techniques similar to the Held-Karp and Gurevich-Shelah methods. These algorithms generalize approaches for the Travelling Salesman Problem and can be applied to a variety of vertex ordering problems, including treewidth and pathwidth. The main contribution lies in providing algorithms with moderately exponential complexity for a broad class of graph ordering problems.
@@ 2.2 Algorithm
The provided algorithm is implemented in a straightforward manner to efficiently solve the ranking problem using the Minimum Weighted Feedback Arc Set (MWFAS) approach. The graph is initially represented using an adjacency list and a weights dictionary, which allows efficient traversal and edge weight management. The process starts by reading the input graph, then cycles are iteratively removed using a depth-first search (DFS) approach to ensure the graph becomes a Directed Acyclic Graph (DAG). We use DFS to effectively check and re-add edges, maintaining connectivity without creating cycles. The use of a heap-based topological sort helps generate a ranking for the vertices, while any ties are broken using the sum of outgoing edge weights. By combining DFS and heap structures, this approach maintains both simplicity and efficiency, offering a practical solution to handle large graphs while minimizing computational overhead.
@@ 3 Experimental Results
Table 1: Number of vertices and edges for each Football dataset.
England_2009_2010 0.13 0.53 0.63 0.01 0.15 0.61 0.46
England_2010_2011 0.23 0.94 0.87 0.04 0.20 0.95 0.65
England_2011_2012 0.18 0.72 0.75 0.04 0.29 0.80 0.53
England_2012_2013 0.17 0.70 0.75 0.04 0.21 0.80 0.51
England_2013_2014 0.13 0.55 0.65 0.03 0.14 0.56 0.46
England_2014_2015 0.19 0.78 0.78 0.05 0.24 0.96 0.69
Average 0.17 0.70 0.74 0.04 0.21 0.78 0.55
Table 2: Loss function values and running time for each Footbal dataset.
Dataset Name Number of Vertices Number of Edges
@@
Table 3: Number of vertices and edges for each basketball dataset.
Basketball_1985 0.14 0.58 0.55 0.97 0.18 0.71 0.46
Basketball_1986 0.15 0.60 0.53 0.87 0.17 0.69 0.42
Basketball_1987 0.16 0.64 0.58 1.21 0.19 0.77 0.48
Basketball_1988 0.15 0.62 0.56 1.21 0.18 0.70 0.45
Basketball_1989 0.15 0.61 0.56 1.25 0.17 0.70 0.46
Basketball_1990 0.14 0.57 0.55 1.00 0.17 0.70 0.44
Basketball_1991 0.15 0.62 0.56 1.19 0.18 0.70 0.45
Basketball_1992 0.14 0.57 0.54 1.09 0.17 0.67 0.43
Basketball_1993 0.14 0.57 0.55 0.86 0.17 0.68 0.44
Basketball_1994 0.13 0.54 0.52 0.86 0.17 0.69 0.42
Basketball_1995 0.14 0.58 0.55 0.87 0.18 0.72 0.44
Basketball_1996 0.16 0.62 0.58 0.90 0.19 0.77 0.48
Basketball_1997 0.16 0.62 0.58 0.91 0.19 0.76 0.48
Basketball_1998 0.14 0.57 0.57 1.00 0.18 0.74 0.46
Basketball_1999 0.16 0.63 0.58 0.91 0.18 0.74 0.48
Basketball_2000 0.15 0.59 0.59 0.99 0.19 0.77 0.50
Basketball_2001 0.14 0.57 0.58 1.03 0.18 0.71 0.47
Basketball_2002 0.15 0.62 0.60 0.97 0.18 0.73 0.48
Basketball_2003 0.16 0.64 0.59 1.02 0.19 0.78 0.51
Basketball_2004 0.15 0.59 0.57 1.03 0.17 0.69 0.46
Basketball_2005 0.15 0.58 0.59 1.05 0.18 0.73 0.49
Basketball_2006 0.15 0.60 0.60 1.10 0.19 0.74 0.49
Basketball_2007 0.15 0.61 0.60 1.16 0.19 0.77 0.49
Basketball_2008 0.16 0.63 0.62 1.33 0.20 0.78 0.51
Basketball_2009 0.15 0.60 0.59 1.18 0.18 0.72 0.47
Basketball_2010 0.15 0.60 0.59 1.26 0.18 0.71 0.49
Basketball_2011 0.16 0.62 0.60 1.27 0.18 0.74 0.49
Basketball_2012 0.14 0.55 0.59 1.53 0.17 0.69 0.48
Basketball_2013 0.15 0.60 0.61 1.27 0.19 0.76 0.51
Basketball_2014 0.16 0.63 0.60 1.28 0.19 0.75 0.49
Table 4: Loss function values and running time for each basketball dataset.
Dataset Name Number of Vertices Number of Edges Naive Loss Simple Loss Ratio Loss Time (Seconds)
Business FM Full 113 1787 0.09 0.37 0.57 0.21
Business FM Assoc 113 1416 0.06 0.26 0.54 0.05
Business FM Asst 113 1345 0.07 0.28 0.60 0.08
Business FM All 113 2796 0.09 0.39 0.55 0.15
Average 113 1836 0.08 0.32 0.55 0.12
Table 5: Summary of Business FM datasets and results.
Dataset Name Number of Vertices Number of Edges Naive Loss Simple Loss Ratio Loss Time (Seconds)
Business F Full 113 427 0.05 0.23 0.61 0.03
Business F All 113 1196 0.09 0.37 0.58 0.06
Business F Assoc 113 500 0.02 0.10 0.54 0.04
Business F Asst 113 612 0.07 0.30 0.61 0.03
Average 113 683.75 0.06 0.25 0.58 0.04
Table 6: Summary of Business F datasets and results.
@@
HeadtoHead 602 5002
Table 7: Number of vertices and edges for the HeadtoHead dataset.
HeadtoHead 0.18 0.75 0.75 4.67 0.24 0.96 0.68
Table 8: Loss function values and running time for the HeadtoHead dataset.
Dataset Name Number of Vertices Number of Edges Naive Loss Simple Loss Ratio Loss Time (Seconds)
ComputerScience_F_All 206 579 0.02 0.08 0.59 0.05
ComputerScience_F_Assoc 206 234 0.00 0.01 0.54 0.04
ComputerScience_F_Asst 206 172 0.00 0.00 0.53 0.06
ComputerScience_F_Full 206 223 0.01 0.05 0.54 0.03
Average 206 302 0.00 0.03 0.55 0.05
Table 9: Summary of Computer Science F datasets and results.
Dataset Name Number of Vertices Number of Edges Naive Loss Simple Loss Ratio Loss Time (Seconds)
History_FM_All 145 2204 0.04 0.19 0.52 0.21
History_FM_Assoc 145 1140 0.02 0.09 0.54 0.07
History_FM_Asst 145 681 0.02 0.08 0.50 0.03
History_FM_Full 145 1204 0.04 0.18 0.58 0.07
Average 145 1307.25 0.03 0.13 0.53 0.09
Table 10: Summary of History FM datasets and results.
@@
We should clarify that in our view, upset naive loss and upset simple loss are not the best functions for qualifying ranking based on weighted comparisons because they do not consider the weight of the violated comparisons; however, we have measured them for our algorithm to compare it with previous works.

### v2

@@ Abstract
The problem of deriving global rankings from pairwise comparisons has received significant attention in the literature. Recently, [He et al. ICML2022] introduced learning-based methods for this task, demonstrating improvements over previous approaches. However, their experimental evaluation did not explore the connections with the closely related Minimum Weighted Feedback Arc Set (MWFAS) problem. This paper shows that simple, learning-free combinatorial algorithms for MWFAS outperform significantly these learning-based methods, achieving both faster runtimes and better ranking quality on the metrics and benchmarks established in [He et al. ICML2022]. [^1] This is a preliminary report.
@@ 1.2.1 Previous Work About Minimum Weighted Feedback Arc Set
  • The Minimum Weighted Feedback Arc Set (MWFAS) problem has received significant attention, particularly regarding approximation algorithms. [23] introduced this problem and proved it is an NP-HARD problem. Early work by [12] provided an approximation algorithm for the MWFAS problem, followed by improvements by [11], that introduced a divide-and-conquer approach using spreading metrics. Although effective, these algorithms used the ellipsoid method for solving linear programming relaxations, making them computationally slow in practice.
@@
    [21] proposed an $O(VE^{4})$ algorithm for MWFAS. However, as the number of edges in the graph increases, the algorithm’s performance becomes slower.
    [29] addressed the parameterized complexity of the feedback vertex and arc set problems in tournaments, focusing on both weighted and unweighted versions. It presents fixed parameter tractable (FPT) algorithms for these problems, including an $O((2.4143)^{k}n^{\omega})$ algorithm for the weighted feedback vertex set problem and an $O((2.415)^{k}n^{\omega})$ algorithm for the weighted feedback arc set problem where $\omega$ is exponent of the running time of the best matrix multiplication algorithm, and $k$ is the size of the feedback set. The paper also introduces algorithms for the parametric duals of these problems and explores their complexity in oriented directed graphs and general directed graphs.
    [19] addressed the problem of finding the largest subset of edges in a directed graph such that the resulting subgraph is acyclic. The authors propose randomized and deterministic approximation algorithms that achieve bounds better than previously known, specifically focusing on graphs without two-cycles and improving performance guarantees for low-degree graphs. The work also introduces methods for handling two-cycles optimally and establishes tight bounds for large classes of graphs.
    [31] explored the use of satisfiability-modulo theories (SMT) solvers to solve the Minimum Feedback Arc Set (MFAS) problem. It presents an SMT encoding for MFAS and highlights the use of incremental solving features, such as unsatisfiable cores. The results show that while the SMT-based approach has potential, it remains computationally infeasible for larger instances, with solving times that scale poorly. The authors also provide complexity bounds, demonstrating that the MFAS problem remains challenging even with advanced SMT techniques.
@@ 1.2.2 Previous Work About Ranking Problem
The ranking problem has also been studied extensively in the literature.[22] introduced a novel ranking method based on pairwise comparisons, aimed at maximizing the probability of full agreement between a ranking and all corresponding pairwise comparisons. This method was demonstrated through portfolio analysis and was compared with previous approaches, highlighting its potential for improved prediction accuracy and indicating areas for further research.[9] introduced Sync-Rank, a method that employs eigenvector and semidefinite programming synchronization to solve ranking, constrained ranking, and rank aggregation problems. The approach leverages spectral or SDP relaxations, followed by rounding procedures to recover rankings. Sync-Rank outperformed other algorithms in both synthetic and real-world datasets while supporting semi-supervised ranking and identifying consistent partial rankings. [8] addressed the problem of ranking players using incomplete pairwise comparison data within the Bradley-Terry-Luce model framework. It established the minimax rate for this ranking task with respect to Kendall’s tau distance, identifying a unique transition between exponential and polynomial rates based on the signal-to-noise ratio. The paper proposed a divide-and-conquer algorithm that achieved the minimax rate, demonstrating its theoretical optimality and computational efficiency.
@@
[17] introduceD a novel pipeline called Graph Neural Re-Ranking (GNRR), which integrates Graph Neural Networks (GNNs) into the re-ranking phase of information retrieval systems. By constructing corpus graphs to model semantic and lexical document relationships, the GNRR approach captures cross-document interactions, improving ranking performance on popular metrics compared to traditional uni-variate baselines. Experimental results show consistent performance gains, especially in challenging datasets, highlighting the potential of GNNs to refine document rankings effectively. The most similar work to ours is [6], which focuses on vertex ordering in directed graph streams. They addressed the Feedback Arc Set problem and propose removing feedback arcs for conducting topological sorts. However, unlike our work, their focus is on unweighted tournament graphs without experimental validation, whereas we address weighted general graphs with extensive experiments.
[1] addresseD the feedback arc set problem and rank aggregation primarily in tournaments, a restricted class of directed graphs where every pair of vertices is connected by a directed edge. In contrast, our work generalizes the problem to arbitrary directed graphs, significantly broadening its applicability. Furthermore, while their study focuses on theoretical approximation algorithms without implementation or empirical validation, we provide practical heuristics accompanied by implementations and extensive experiments, demonstrating the effectiveness of our approach on real-world datasets.
[5] presented exact algorithms for solving several vertex ordering problems on graphs. The proposed algorithms operate in either $O^{*}(2^{n})$ time and space or $O^{*}(4^{n})$ time and polynomial space, leveraging dynamic programming and recursive techniques similar to the Held-Karp and Gurevich-Shelah methods. These algorithms generalize approaches for the Travelling Salesman Problem and can be applied to a variety of vertex ordering problems, including treewidth and pathwidth. The main contribution lies in providing algorithms with moderately exponential complexity for a broad class of graph ordering problems.
[7] addressed the problem of fair rank aggregation by introducing algorithms that ensure fairness across diverse groups in ranked outputs. The authors propose exact and approximation algorithms for the Closest Fair Ranking and Fair Rank Aggregation problems under various fairness notions, leveraging metrics such as Kendall tau and Ulam distances to achieve efficient and principled solutions.
@@ 2.2 Algorithm
Algorithm 1 is implemented in a straightforward manner to efficiently solve the ranking problem using the Minimum Weighted Feedback Arc Set (MWFAS) approach. The graph is initially represented using an adjacency list and a weights dictionary, which allows efficient traversal and edge weight management. The process starts by reading the input graph, then cycles are iteratively removed using a depth-first search (DFS) approach to ensure the graph becomes a Directed Acyclic Graph (DAG). We use DFS to effectively check and re-add edges, maintaining connectivity without creating cycles. The use of a heap-based topological sort helps generate a ranking for the vertices, while any ties are broken using the sum of outgoing edge weights. By combining DFS and heap structures, this approach maintains both simplicity and efficiency, offering a practical solution to handle large graphs while minimizing computational overhead.
@@ 3 Experimental Results
England_2009_2010 0.13 0.53 0.60 0.01 0.15 0.61 0.46
England_2010_2011 0.23 0.94 0.84 0.04 0.20 0.95 0.65
England_2011_2012 0.18 0.72 0.73 0.04 0.29 0.80 0.53
England_2012_2013 0.17 0.70 0.73 0.04 0.21 0.80 0.51
England_2013_2014 0.13 0.55 0.62 0.03 0.14 0.56 0.46
England_2014_2015 0.19 0.78 0.78 0.05 0.24 0.96 0.69
Business FM Full 0.09 0.37 0.48 0.30 0.09 0.36 0.31
computerScience_FM_Full 0.05 0.21 0.51 0.54 0.08 0.32 0.26
History_FM_Full 0.04 0.18 0.49 0.73 0.07 0.28 0.21
Basketball_1985 0.11 0.46 0.53 1.20 0.18 0.71 0.46
Basketball_1986 0.12 0.50 0.52 1.31 0.17 0.69 0.42
Basketball_1987 0.13 0.52 0.56 2.20 0.19 0.77 0.48
Basketball_1988 0.12 0.48 0.52 1.82 0.18 0.70 0.45
Basketball_1989 0.13 0.52 0.54 1.49 0.17 0.70 0.46
Basketball_1990 0.12 0.48 0.52 1.34 0.17 0.70 0.44
Basketball_1991 0.13 0.53 0.54 1.35 0.18 0.70 0.45
Basketball_1992 0.11 0.46 0.52 1.37 0.17 0.67 0.43
Basketball_1993 0.12 0.48 0.54 1.33 0.17 0.68 0.44
Basketball_1994 0.11 0.47 0.52 1.41 0.17 0.69 0.42
Basketball_1995 0.11 0.47 0.52 1.55 0.18 0.72 0.44
Basketball_1996 0.13 0.55 0.55 1.63 0.19 0.77 0.48
Basketball_1997 0.12 0.51 0.57 1.76 0.19 0.76 0.48
Basketball_1998 0.12 0.48 0.56 1.83 0.18 0.74 0.46
Basketball_1999 0.13 0.53 0.56 2.39 0.18 0.74 0.48
Basketball_2000 0.13 0.55 0.57 2.29 0.19 0.77 0.50
Basketball_2001 0.12 0.50 0.56 2.03 0.18 0.71 0.47
Basketball_2002 0.13 0.54 0.58 1.78 0.18 0.73 0.48
Basketball_2003 0.15 0.60 0.60 1.96 0.19 0.78 0.51
Basketball_2004 0.12 0.49 0.55 1.80 0.17 0.69 0.46
Basketball_2005 0.13 0.53 0.57 2.16 0.18 0.73 0.49
Basketball_2006 0.13 0.54 0.58 2.60 0.19 0.74 0.49
Basketball_2007 0.14 0.56 0.58 2.96 0.19 0.77 0.49
Basketball_2008 0.14 0.57 0.61 2.39 0.20 0.78 0.51
Basketball_2009 0.13 0.52 0.58 2.85 0.18 0.72 0.47
Basketball_2010 0.13 0.53 0.58 2.47 0.18 0.71 0.49
Basketball_2011 0.14 0.56 0.59 2.42 0.18 0.74 0.49
Basketball_2012 0.13 0.52 0.57 2.61 0.17 0.69 0.48
Basketball_2013 0.14 0.57 0.59 2.79 0.19 0.76 0.51
Basketball_2014 0.15 0.69 0.63 3.05 0.20 0.78 0.49
Basketball_finer1985 0.14 0.55 0.13 3.61 0.18 0.71 0.00
Basketball_finer1986 0.15 0.59 0.13 2.76 0.17 0.69 0.00
Basketball_finer1987 0.15 0.60 0.13 2.77 0.19 0.77 0.00
Basketball_finer1988 0.15 0.59 0.13 2.87 0.18 0.70 0.00
Basketball_finer1989 0.15 0.58 0.13 3.54 0.18 0.70 0.00
Basketball_finer1990 0.15 0.58 0.13 4.42 0.18 0.71 0.00
Basketball_finer1991 0.15 0.61 0.13 3.50 0.18 0.71 0.00
Basketball_finer1992 0.14 0.55 0.13 3.55 0.17 0.67 0.00
Basketball_finer1993 0.14 0.57 0.13 3.41 0.17 0.68 0.00
Basketball_finer1994 0.14 0.55 0.14 3.16 0.17 0.67 0.00
Basketball_finer1995 0.14 0.54 0.13 3.49 0.18 0.72 0.00
Basketball_finer1996 0.16 0.63 0.12 3.86 0.19 0.77 0.01
Basketball_finer1997 0.15 0.62 0.13 3.69 0.19 0.75 0.01
Basketball_finer1998 0.14 0.57 0.13 3.87 0.18 0.73 0.01
Basketball_finer1999 0.16 0.62 0.12 4.10 0.18 0.73 0.01
Basketball_finer2000 0.16 0.65 0.13 4.77 0.19 0.78 0.01
Basketball_finer2001 0.16 0.63 0.13 4.74 0.18 0.73 0.01
Basketball_finer2002 0.17 0.69 0.13 5.10 0.19 0.77 0.01
Basketball_finer2003 0.15 0.60 0.12 5.31 0.19 0.78 0.01
Basketball_finer2004 0.16 0.64 0.13 4.89 0.18 0.71 0.01
Basketball_finer2005 0.17 0.66 0.13 7.23 0.19 0.75 0.01
Basketball_finer2006 0.17 0.68 0.14 5.97 0.19 0.76 0.01
Basketball_finer2007 0.17 0.68 0.14 6.80 0.20 0.80 0.01
Basketball_finer2008 0.16 0.65 0.14 6.67 0.20 0.78 0.01
Basketball_finer2009 0.16 0.62 0.14 9.65 0.19 0.75 0.01
Basketball_finer2010 0.17 0.67 0.14 8.06 0.19 0.75 0.01
Basketball_finer2011 0.17 0.62 0.14 6.87 0.19 0.77 0.01
Basketball_finer2012 0.15 0.67 0.14 7.20 0.18 0.74 0.01
Basketball_finer2013 0.17 0.69 0.13 7.30 0.20 0.78 0.01
Basketball_finer2014 0.15 0.69 0.13 8.36 0.20 0.78 0.00
HeadtoHead 0.18 0.73 0.75 5.92 0.24 0.96 0.68
Animal 0.07 0.30 0.38 0.26 0.06 0.25 0.22
Football_finer(2009) 0.20 0.78 0.31 0.40 0.16 0.65 0.17
Football_finer(2010) 0.28 1.12 0.29 0.20 0.29 0.99 0.17
Football_finer(2011) 0.21 0.82 0.29 0.07 0.21 0.84 0.17
Football_finer(2012) 0.20 0.81 0.26 0.05 0.21 0.86 0.15
Football_finer(2013) 0.16 0.63 0.28 0.05 0.14 0.57 0.19
Football_finer(2014) 0.26 1.05 0.47 0.03 0.26 1.00 0.34
Table 1: Loss function values and running times for all datasets.
@@
Business FM Full 113 1787
computerScience_FM_Full 206 1407
History_FM_Full 145 1204
@@
Table 2: Number of vertices and edges for various datasets (Part 1).
Dataset Name Number of Vertices Number of Edges
@@
HeadtoHead 602 5002
Animal 19 193
Basketball_finer1985 282 4814
Basketball_finer1986 283 4862
Basketball_finer1987 290 5088
Basketball_finer1988 290 5170
Basketball_finer1989 293 5318
Basketball_finer1990 292 5350
Basketball_finer1991 295 5420
Basketball_finer1992 298 5444
Basketball_finer1993 298 5160
Basketball_finer1994 301 5252
Table 3: Number of vertices and edges for various datasets (Part 2).
@@
Basketball_finer1995 302 5336
Basketball_finer1996 305 5498
Basketball_finer1997 305 5628
Basketball_finer1998 306 5684
Basketball_finer1999 310 5788
Basketball_finer2000 318 6274
Basketball_finer2001 318 6116
Basketball_finer2002 321 6192
Basketball_finer2003 327 6356
Basketball_finer2004 326 6316
Basketball_finer2005 330 6476
Basketball_finer2006 334 6680
Basketball_finer2007 336 7186
Basketball_finer2008 342 7386
Basketball_finer2009 347 7478
Basketball_finer2010 347 7538
Basketball_finer2011 345 7504
Basketball_finer2012 345 7580
Basketball_finer2013 347 7616
Basketball_finer2014 351 7650
Football_finer(2009) 20 380
Football_finer(2010) 20 380
Football_finer(2011) 20 380
Football_finer(2012) 20 380
Football_finer(2013) 20 380
Football_finer(2014) 20 300
Table 4: Number of vertices and edges for Basketball Finer and Football Finer datasets.
@@
Input: $adjacency\_matrix$, $scores$, $index$, $lower\_bound$, $upper\_bound$, $epsilon$, $steps$
Output: Optimal score value for $index$
for $step\leftarrow 1$ to $steps$ do
       $mid1\leftarrow lower\_bound+\frac{upper\_bound-lower\_bound}{3}$;
       $mid2\leftarrow upper\_bound-\frac{upper\_bound-lower\_bound}{3}$;
       $scores[index]\leftarrow mid1$;
       $loss1\leftarrow\text{ComputeRatioUpsetLoss}(adjacency\_matrix,scores)$;
       $scores[index]\leftarrow mid2$;
       $loss2\leftarrow\text{ComputeRatioUpsetLoss}(adjacency\_matrix,scores)$;
       if $loss1<loss2$ then
             $upper\_bound\leftarrow mid2$;
       else
             $lower\_bound\leftarrow mid1$;
      if $upper\_bound-lower\_bound<epsilon$ then
             break;
$scores[index]\leftarrow\frac{lower\_bound+upper\_bound}{2}$;
return $scores[index]$;
Algorithm 2 Trinary Search Optimization for Minimizing Ratio Upset Loss
Input: $adjacency\_matrix$, $scores$,$num\_iterations$
Output: Optimized $scores$
for $iteration\leftarrow 1$ to $num\_iterations$ do
       $sorted\_indices\leftarrow\text{argsort}(scores)$;
        /* Sort indices based on scores */
       for $i\leftarrow 1$ to length$(sorted\_indices)$ do
             $index\leftarrow sorted\_indices[i]$;
             if $i=1$ then
                   $lower\_bound\leftarrow 0$;
                   $upper\_bound\leftarrow scores[sorted\_indices[i+1]]$;
                    /* First element bounds */
             else if $i=\text{length}(sorted\_indices)$ then
                   $lower\_bound\leftarrow scores[sorted\_indices[i-1]]$;
                   $upper\_bound\leftarrow\max(scores)+1$;
                    /* Last element bounds */
             else
                   $lower\_bound\leftarrow scores[sorted\_indices[i-1]]$;
                   $upper\_bound\leftarrow scores[sorted\_indices[i+1]]$;
                    /* Middle element bounds */
            $scores[index]\leftarrow\textsc{TrinarySearchOptimize}(adjacency\_matrix,scores,index,lower\_bound,upper\_bound,\epsilon,\text{steps})$;
              /* Trinary search optimization */
return $scores$;
Algorithm 3 Minimize Ratio Upset Loss
3.1 Trying to decrease ratio loss without increasing naive loss or simple loss
In experiments, we observed that while our algorithm’s naive loss and simple loss are better than the state-of-the-art, it is not the case for the ratio loss. Therefore, we designed an algorithm to decrease the ratio loss without changing the ranking order. We should note that the ratio loss depends not only on the ranking of the elements but also on the scores we assign to the elements. Therefore, we designed Algorithm 3 to update the scores of the elements without changing the rankings. Without loss of generality, we can consider that $scores_{1}\leq scores_{2}\leq...\leq scores_{n}$. In each algorithm step, it iterates over all of the scores. For element $i$, the algorithm does a trinary search between $scores_{i-1}$ and $scores_{i+1}$ to find the optimal amount for $scores_{i}$. Please note that because still $scores_{i}$ will remain between $scores_{i-1}$ and $scores_{i+1}$, naive loss and simple loss do not change, but ratio loss may be decreased. The reason that trinary search works well is that based on our experiments, with a very good approximation, we can consider that ratio loss function for $scores_{i}$ from $scores_{i-1}$ to $scores_{i+1}$, when other elements of $scores$ array do not change, is a unimodal function. The results of using  Algorithm 2 for 40 iterations can be seen for some datasets in  Table 5 It is crucial to note that maybe no solution exists that is better in all three loss functions than the GNNRank algorithm.
Dataset Name Initial Main Algorithm Ratio Loss GNNRank Ratio Loss Main Algorithm Ratio Loss After 40 rounds of Optimization
England_2009_2010 0.6 0.46 0.52
England_2010_2011 0.84 0.65 0.78
England_2011_2012 0.73 0.53 0.63
England_2012_2013 0.73 0.51 0.6
England_2013_2014 0.62 0.46 0.52
England_2014_2015 0.78 0.69 0.74
Average 0.72 0.55 0.63
Business FM Full 0.48 0.31 0.44
computerScience_FM_Full 0.51 0.26 0.45
History_FM_Full 0.49 0.21 0.4
Basketball_1985 0.53 0.46 0.51
Basketball_1986 0.52 0.42 0.5
Basketball_1987 0.56 0.48 0.54
Basketball_1988 0.52 0.45 0.51
Basketball_1989 0.54 0.46 0.52
Basketball_1990 0.52 0.44 0.51
Basketball_1991 0.54 0.45 0.52
Basketball_1992 0.52 0.43 0.5
Basketball_1993 0.54 0.44 0.52
Basketball_1994 0.52 0.42 0.51
Basketball_1995 0.52 0.44 0.52
Table 5: Comparison of Ratio Loss, GNNRank Ratio Loss, and Main Algorithm Ratio Loss (40 rounds)
3.2 Analysing the results